OpenStudy (erniellyd28):

4- 3 5 ---- 8 ------------ 1 3 2 ---- - ---- 2 4

OpenStudy (anonymous):

No problem!!! Okay, the problem 4 - 3 5/8 ---------- 2 1/2 - 3/4 So to get the answer you must first do; 2 1/2 - 3/4 change 2 1/2 to 5/2 so now you have 5/2 - 3/4 which is 7/4 so know you have 4 - 3 5/8 ---------- 7/4 Now lets do the top change 3 5/8 to 29/8 so know its 4- 29/8 which is 3/8 so know 3/8 ----- 7/4 now remember how to divide do; 3/8 X 4/7 which is 3/14 and thats the answer, 3/14! Helped?

OpenStudy (anonymous):

\[ \frac{4-3\frac58}{2\frac12-\frac34}=\frac{4-\frac{29}8}{\frac52-\frac34}=\frac{32-29}{20-6}=\frac{3}{14} \]

OpenStudy (anonymous):

Yeh just message me them like before;)